# BED-LLM: Intelligent Information Gathering with LLMs and Bayesian Experimental Design

### v2

@@ 3.1 Generating Candidate Questions
As it is not computationally feasible to directly optimize over the space of possible questions, we rely on using the LLM to propose diverse candidate questions, $\mathcal{X}^{\text{cand}}$, then select the best question from these. We consider two specific approaches: 1) Unconstrained generation. Given $h_{t-1}$, the LLM is simply asked to propose new questions by sampling from $p_{\mathrm{LLM}}(x_{t};h_{t-1})$ with appropriate prompting. 2) Conditional generation. The LLM is given both $h_{t-1}$ and a generated set of hypotheses $\Theta^{\text{cand}}=\{\theta^{(n)}\}_{n=1}^{N}$, such that we sample from $p_{\mathrm{LLM}}(x_{t};[h_{t-1},\Theta^{\text{cand}}])$. Specifically, the LLM is prompted to propose questions that “slice” the hypothesis pool into roughly balanced subsets.
For both strategies, we sample $M$ questions jointly with a relatively high temperature to encourage diversity. Conditional generation allows us to “guide” the LLM to propose highly informative questions. However, it risks overfitting to $\Theta^{\text{cand}}$. In practice, we find it is effective for discrete spaces (§6.1), but less so for spaces with complex, overlapping hypotheses (§6.2). We restrict questions to multiple-choice format to allow $p_{\mathrm{LLM}}\left(y_{t};\left[\theta,x_{t}\right]\right)$ to produce well-calibrated probabilities (see §4).
@@ 3.2 Estimating EIG for each question
To estimate the EIG based on Equation 2 for a given question $x_{t}$, we derive the following Rao-Blackwellized estimator based on the LLM’s predictive distribution:
$\begin{split}\text{EIG}_{\theta}(x_{t};h_{t-1})\approx&\textstyle\frac{1}{N}\sum_{n=1}^{N}\sum_{y_{t}\in\mathcal{Y}}p_{\mathrm{LLM}}(y_{t};[\theta^{(n)},x_{t}])\log p_{\mathrm{LLM}}(y_{t};[\theta^{(n)},x_{t}])\\ &-\textstyle\sum_{y_{t}\in\mathcal{Y}}\hat{p}(y_{t};[h_{t-1},x_{t}])\log\hat{p}(y_{t};[h_{t-1},x_{t}]),\end{split}$ (3)
where $\hat{p}(y_{t};[h_{t-1},x_{t}])\!:=\!\frac{1}{N}\sum_{n=1}^{N}p_{\mathrm{LLM}}(y_{t};[\theta^{(n)},x_{t}])$ and $\theta^{(n)}\!\sim\!p_{f}(\theta;h_{t-1})$, see §3.3. This estimator has been used in other BED contexts (Gal et al., 2017; Rainforth, 2017). Note that the samples do not need to be independent for this estimator to converge, provided they satisfy some appropriate form of ergodicity or decaying correlation (see, e.g., Billingsley (2013)). When constructing this estimator, we compute the $p_{\mathrm{LLM}}(y_{t};[\theta^{(n)},x_{t}])$ terms using the LLM’s logits whenever possible. By the Rao-Blackwell theorem, this always produces lower variance than purely sample–based estimators (Rao et al., 1945), like those employed in Hu et al. (2024) and Kobalczyk et al. (2025).
@@
This is problematic as, in practice, the expected likelihood entropy will vary with $x_{t}$. For example, if $\theta=$“Dog”, then the response to the question “Is it an animal?” should be close to deterministic, but the expected response to the question “Does it have black fur?” clearly is not. In general, $\mathbb{E}_{p(\theta;h_{t-1})}[\mathrm{H}[p(y_{t}|\theta;x_{t},h_{t-1})]]$ measures how certainly the question can be answered once $\theta$ is known. Including it in our objective is essential in avoiding questions that are irrelevant, ambiguous, unclear, or simply unhelpful in our quest to learn about $\theta$. We provide an illustrative example of this in Section˜A.2.1 and empirical evidence for this in §6. Given that approximating the EIG with marginal predictive entropy does not provide meaningful computational savings (as it does not reduce the required number of LLM calls), we advise against making such deterministic likelihood assumptions.
3.3 Prior Construction and Belief Updating
The Savage axioms (Savage, 1954) tell us that a rational agent should update its beliefs in a Bayesian manner. However, doing full Bayesian updates to our model as the history grows is generally impractical for computational reasons in the LLM setting, as it requires approximate inference and this, in turn, typically requires large numbers of expensive likelihood evaluations. Furthermore, the Savage axioms only hold if our (implied) prior truly represents our beliefs, but we find that $p_{\mathrm{LLM}}(\theta)$ is typically heavily overconfident on a small number of possible hypotheses and can struggle to convey the full range of possibilities even with careful prompting and a high temperature (c.f. Figure˜5).
To avoid these shortfalls, we instead propose an approach that balances tractability and faithfulness. Although we will still use $p_{\mathrm{LLM}}(\theta;h_{t-1})$ as the basis for deriving our belief state over $\theta$ (i.e. our intermediate prior), we make various alterations to effectively incorporate historical information and ensure diversity. Our derived distribution, which we refer to as $p_{f}(\theta;h_{t-1})$, differs from $p_{\mathrm{LLM}}(\theta;h_{t-1})$ in two key ways. First, we filter the generated hypotheses according to whether they are compatible with the history $h_{t-1}$. We do this by using the LLM to zero-shot check the compatibility of each sampled $\theta$ with all the previous question–answer pairs in $h_{t-1}$ (using $p_{\mathrm{LLM}}(y_{i};[\theta,x_{i}])\forall i=1:t-1$) and then rejecting that sample if an incompatibility is found. Specifically, a sample is rejected if the likelihood of an observed answer falls below a predefined threshold, chosen to balance robustness to model uncertainty against the need to enforce strict historical coherence. To reduce the computational cost of generating and evaluating hypotheses, we further include a hypothesis-retention mechanism: any hypotheses from the previous turn which remain consistent with the most recent question and observation are retained in the hypothesis set without regeneration. Second, we make a number of modifications to promote diversity. Rather than generate candidates independently, we prompt the LLM to generate batches of candidates using a prompt encouraging diversity. After filtering these candidates as above and removing duplicates, we then impose a uniform distribution. Details of our exact setup for $p_{f}(\theta;h_{t-1})$ are given in Appendix˜E.
@@ 4 On the Specification of $p(\theta,y_{t};h_{t-1},x_{t})$, and its Implications
As we described in §3, successfully applying sequential BED in the LLM setting hinges upon how we specify, and update, the joint distribution $p(\theta,y_{t};h_{t-1},x_{t})$. In particular, as previously highlighted, there are two distinct ways to derive the joint model from our LLM: using a prior–likelihood pairing, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p(\theta;h_{t-1})p\left(y_{t};\left[\theta,x_{t}\right]\right)$}, or a data–estimation pairing, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t};[h_{t-1},x_{t}]\right)p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$}. The first construction mirrors deriving our beliefs about $\theta$ from a conventional Bayesian posterior with a concrete prior and likelihood derived (at least partially) from the LLM, whereas the second has analogies to a marginal-posterior approach (Fong et al., 2023; Falck et al., 2024) in that it that samples hypothetical data and draws inferences on $\theta$ given hypothetical data using in-context learning. In our outlined BED-LLM approach, we adopted a prior–likelihood pairing. Below, we justify this decision and also discuss certain settings where the data–estimation setup might be preferable instead.
Modeling flexibility   The most obvious relative merits of the prior–likelihood and data–estimation pairings are in the flexibility in how each term is chosen. The prior–likelihood pairing gives us greater flexibility to construct a prior set of beliefs over $\theta$ that is distinct to the LLM’s internal beliefs, as it allows us to directly control this prior by changing $p(\theta;h_{t-1})$, whereas the prior is only implicitly defined in the data–estimation pairing. In §3.3 we exploited this flexibility through our definition of $p_{f}(\theta;h_{t-1})$. On the other hand, the data–estimation pairing could provide some beneficial flexibility in specifying how the data itself is simulated through changing $p\left(y_{t};[h_{t-1},x_{t}]\right)$, which could, for example, be useful when we have access to external data simulators.
@@
This difference becomes particularly noticeable when the complexities of the spaces of $\theta$ and $y$ differ significantly. Our ability to draw sensible samples of either will generally be quite robust to these spaces being complex or high–dimensional; this is where LLMs tend to thrive, effectively generating highly complex outputs in an autoregressive manner. However, evaluating the entropy of a distribution becomes dramatically harder as the dimensionality or complexity increases (Acharya et al., 2019; Paninski, 2003), and the entropy of the predictive distribution of an LLM in such cases will not typically provide a sensible measure of uncertainty (Kadavath et al., 2022; Desai & Durrett, 2020). As such, the decision between joint formulations should predominantly be based on the complexity of the space of $\theta$ versus that of $y$: we should generally favor the prior–likelihood formulation if $\theta$ is more complex and the data–estimation formulation if $y$ is more complex. For the problems that we consider, the space of $y$ is less complex than that of $\theta$, indicating we should, in general, use the prior–likelihood formulation. However, in cases where this is not true, the data–estimation formulation may be preferable instead. We provide additional discussion on the impact of the chosen pairing on our entropy estimate, plus discussion on the choice of $\theta$, in Appendix˜B.
@@ 5 Related Work
Several works have explored the baseline ability of LLMs to rapidly learn about a parameter of interest by asking questions (Zhang et al., 2024; Li et al., 2025b)—effectively our Naive baseline in §6. While these works demonstrate some ability to adaptively construct information-seeking questions, they often fail to extract important information (Li et al., 2025a).
Some works have further specifically attempted to choose questions based on model-based uncertainty criteria (Piriyakulkij et al., 2023; Hu et al., 2024; Kobalczyk et al., 2025; Mazzaccara et al., 2024; Cooper et al., 2025). None of these works provide the same careful consideration of how the underlying joint model should be formulated, which underpins our own work, and they all assume deterministic likelihood models that mean their objectives correspond to a sample-based estimate of marginal predictive entropy in practice, as explained in Section˜3.2. In general, these previous works have also required restrictions on the space of allowable hypotheses, $\theta$, and typically require additional assumptions and/or approximations. More extensive discussion of related work is given in Appendix˜C.
@@ 6 Experiments
We produce answers to the questioner LLM’s questions using a separate answerer LLM. The answerer is provided with a ground-truth $\theta^{*}$ (a target entity in 20 Questions or a user profile in preference elicitation) and processes individual questions from the questioner without access to any of the questioner’s context (i.e. $h_{t-1}$and the questioner’s prompts). We test two questioner-answerer setups, where the two are served by separate instances of the same LLM, or two different LLMs. The latter scenario is important because in practice, the answerer will often follow a different distribution than the questioner’s internal model for reasoning about responses, thereby forming a model misspecification.
@@
We compare BED-LLM against two existing baseline methods: Naive and Split. Naive involves prompting the questioner to directly generate an informative next question, without explicit hypotheses generation or computing a data-acquisition objective, and then sampling the question with temperature T=1T\!=\!1; this was explored by Zhang et al. (2024). Split involves choosing the question that most equally splits a sampled set of hypotheses $\Theta^{\text{cand}}$, which corresponds to maximizing the marginal predictive entropy $\mathrm{H}[p(y_{t};x_{t},h_{t-1})]$ in a model with a deterministic likelihood. As such, the methods of Cooper et al. (2025), Hu et al. (2024), Kobalczyk et al. (2025), Mazzaccara et al. (2024) and Piriyakulkij et al. (2023) can all be viewed as variants of this Split baseline. While Split is not applicable to the preference-elicitation scenario, where a deterministic likelihood is not viable, it is to our knowledge the state-of-the-art method for 20 Questions. We note that our own Split baseline implementation achieves dramatically better results than reported by, for example, Kobalczyk et al. (2025), so this constitutes a very strong baseline relative to previous work. On top of this, our implementation of Naive appears to signficantly improve over that of Zhang et al. (2024),
@@
We consider three sets of 20 Questions problems: Animals, Celebrities, and Things (See Section˜F.1). Each problem set comprises 100 target entities $\{\theta^{*}_{i}\}_{i=1}^{100}$ from a given category. The space of possible $\theta$ is large and not explicitly defined or restricted: we do not tell the LLM this set of target entities, so the space of $\theta$ is bounded only by what the LLM can generate. We note that by comparison, many previous works have relied on restricted spaces for $\theta$ (Chan et al., 2025; Hu et al., 2024; Piriyakulkij et al., 2023; Wang et al., 2025).
@@
Dataset Model Naive Split BED-LLM Entropy Data–Est. ICL Beliefs Impl. Max.
@@
Table 1: Success rate (%) for 20 Questions at the end of the game. Best result in bold. $\pm$ numbers show the standard error of the mean estimated using $\sqrt{p(1-p)/(n-1)}$ where $p$ is the success percentage and $n$ is the number of datapoints. This estimator is positively biased and thus conservative. Data–Est. is not possible to run for GPT models due to limited logprobs support in OpenAI API.
@@
BED-LLM improves over Naive and Split baselines
Our results in Table˜1 and Figure˜1 show BED-LLM significantly outperforming Naive and Split across all problems and LLMs. Particularly notable is that BED-LLM’s final success rate is typically more than double that of Naive, highlighting the big gains that can be achieved by using explicit EIG maximisation instead of implicit LLM reasoning.
@@
In order to understand the importance of BED-LLM’s algorithmic components, we further evaluate four ablations on this problem: Entropy, Data–Estimation, ICL Beliefs and Implicit Maximization. Each of these differs from BED-LLM with respect to one algorithmic component. Entropy replaces the EIG data-acquisition objective with the marginal predictive entropy $\mathrm{H}[p(y_{t};x_{t},h_{t-1})]$ (Section˜3.2); this contrasts with Split in that it uses $p_{\mathrm{LLM}}(y_{t+1};[\theta,x_{t+1}])$ rather than a deterministic likelihood. Implicit Maximization involves sampling candidate questions and prompting the questioner to select the most informative one, without any explicit objective estimation (Section˜3.2). ICL Beliefs uses $\theta$ belief updates derived from simple in-context learning, namely $p_{\mathrm{LLM}}(\theta;h_{t-1})$ instead of $p_{f}(\theta;h_{t-1})$, testing the importance of our filtering mechanism (Section˜3.3). Data–Estimation uses a data–estimation pairing rather than BED-LLM’s prior–likelihood setup (see Appendix˜D). In Table˜1 we see that BED-LLM comfortably outperforms all alternative approaches. Notably, Entropy provides the strongest ablation baseline, with performance almost identical (sometimes superior) to Split. This shows that the use of a non-deterministic likelihood is beneficial in allowing us to target a proper EIG, rather than because it is particularly detrimental to the marginal predictive entropy itself.
@@
Our analysis in §4 is validated by our results: BED-LLM’s prior-likelihood approach substantially outperforms Data–Estimation. Data–Estimation still outperforms Naive, but interestingly it performs worse than Entropy, highlighting the importance of estimating uncertainty in the $y$ space instead of $\theta$ space. These findings reinforce our claim that the choice of joint-model factorization is a critical algorithmic decision.
@@
We also see the importance of two other aspects of BED-LLM. First, how we produce our beliefs over $\theta$ matters: deriving beliefs using simple in-context learning, as in ICL Beliefs, lead to massive performance drops. Second, while BED-LLM’s routines for sampling candidate questions and hypotheses are crucial, they alone are not sufficient: passing the samples to an LLM and prompting it to select the highest-EIG question, as in Implicit Maximization, works much less well than using the samples to explicitly maximize EIG.
@@
Our results in Figure˜1 demonstrate that the benefit of BED-LLM persists even under model misspecification. This is important for applicability to real-world users, whose responses will follow a different distribution than the questioner LLM.

### v3

@@ 3.2 Generating candidate questions
As it is not computationally feasible to directly optimize over the space of possible questions, we rely on using the LLM to propose diverse candidate questions, $\mathcal{X}^{\mathrm{cand}}$, then select the best question from these. We consider two specific approaches. (a) Unconstrained generation: given $h_{t-1}$, the LLM is simply asked to propose new questions by sampling from $p_{\mathrm{LLM}}(x_{t};h_{t-1})$ with appropriate prompting. (b) Conditional generation: the LLM is given both $h_{t-1}$ and a generated set of hypotheses, $\Theta^{\mathrm{cand}}=\{\theta_{n}\}_{n=1}^{N}$, such that we sample from $p_{\mathrm{LLM}}(x_{t};[h_{t-1},\Theta^{\mathrm{cand}}])$; specifically, the LLM is prompted for questions that “slice” the hypothesis pool into roughly balanced subsets.
For both strategies, we sample $M$ questions jointly with a relatively high temperature to encourage diversity. Conditional generation allows us to “guide” the LLM to propose highly informative questions. However, it risks overfitting to $\Theta^{\mathrm{cand}}$. In practice, we find it is effective for discrete spaces (Section˜6.1), but less so for spaces with complex, overlapping hypotheses (Section˜6.2). We restrict questions to multiple-choice format to simplify uncertainty quantification (see Section˜4).
@@ 3.3 Estimating EIG for each question
To estimate the EIG based on Equation 2 for a given question $x_{t}$, we use the following Rao-Blackwellized estimator based on the LLM’s predictive distribution:
$\begin{split}\mathrm{EIG}_{\theta}(x_{t};h_{t-1})\approx&\textstyle\frac{1}{N}\sum_{n=1}^{N}\sum_{y_{t}\in\mathcal{Y}}p_{\mathrm{LLM}}(y_{t};[\theta_{n},x_{t}])\log p_{\mathrm{LLM}}(y_{t};[\theta_{n},x_{t}])\\ &-\textstyle\sum_{y_{t}\in\mathcal{Y}}\hat{p}(y_{t};[h_{t-1},x_{t}])\log\hat{p}(y_{t};[h_{t-1},x_{t}]),\end{split}$ (3)
where $\hat{p}(y_{t};[h_{t-1},x_{t}])\!:=\!\frac{1}{N}\sum_{n=1}^{N}p_{\mathrm{LLM}}(y_{t};[\theta_{n},x_{t}])$ and $\theta_{n}\!\sim\!p_{f}(\theta;h_{t-1})$ (see Section˜3.1). This estimator has been used in other BED contexts (Gal et al., 2017; Rainforth, 2017). Note that the samples do not need to be independent for this estimator to converge, provided they satisfy some appropriate form of ergodicity or decaying correlation (see, e.g., Billingsley (2013)). When constructing this estimator, we compute the $p_{\mathrm{LLM}}(y_{t};[\theta_{n},x_{t}])$ terms using the LLM’s logits whenever possible. By the Rao-Blackwell theorem, this always produces lower variance than purely sample-based estimators (Rao et al., 1945), like those employed in Hu et al. (2024) and Kobalczyk et al. (2025).
@@
This is problematic as, in practice, the expected likelihood entropy will vary with $x_{t}$. In general, $\mathbb{E}_{p(\theta;h_{t-1})}[\mathrm{H}[p(y_{t}|\theta;x_{t},h_{t-1})]]$ measures how certainly the question can be answered once $\theta$ is known. Including it in our objective is essential in avoiding questions that are irrelevant, ambiguous, unclear, or simply unhelpful in our quest to learn about $\theta$. Figure˜1 illustrates this concretely—question $\tilde{x}_{4}$ has high marginal entropy and low conditional entropy (answers are crisp given $\theta$), yielding high EIG, while $\tilde{x}_{5}$ has equally high marginal entropy but also high conditional entropy (answers are noisy even given $\theta$), so the terms cancel and nearly nothing is learned. An entropy-only method would score both similarly and risk wasting a turn; the same failure mode causes the Entropy ablation to provide no improvement over Prompt-Only in our preference elicitation experiments (Section˜6.2). Given that both terms of Equation˜3 are computed from the same LLM likelihood evaluations, retaining the full EIG requires no additional calls and provides no computational savings. Thus, we advise against making deterministic likelihood assumptions.
@@ 4 On the specification of $p(\theta,y_{t};h_{t-1},x_{t})$, and its implications
As we described in §3, successfully applying sequential BED in the LLM setting hinges upon how we specify, and update, the joint distribution, $p(\theta,y_{t};h_{t-1},x_{t})$. In particular, as previously highlighted, there are two distinct ways to derive the joint model from our LLM: using a prior–likelihood pairing, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\!\left(\theta;h_{t-1}\right)p\!\left(y_{t};\left[\theta,x_{t}\right]\right)$}, or a data–estimation pairing, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\!\left(y_{t};[h_{t-1},x_{t}]\right)p\!\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$}. The first construction mirrors deriving our beliefs about $\theta$ from a conventional Bayesian posterior with a concrete prior and likelihood derived (at least partially) from the LLM, whereas the second has analogies to a marginal-posterior approach (Fong et al., 2023; Falck et al., 2024) in that it that samples hypothetical data and draws inferences on $\theta$ given hypothetical data using in-context learning. In our outlined BED-LLM approach, we adopted a prior–likelihood pairing. Below, we justify this decision and also discuss certain settings where the data–estimation setup might be preferable instead.
Modeling flexibility   The most obvious relative merits of the prior–likelihood and data–estimation pairings are in the flexibility in how each term is chosen. The prior–likelihood pairing gives us greater flexibility to construct a prior set of beliefs over $\theta$ that is distinct to the LLM’s internal beliefs, as it allows us to directly control this prior by changing $p\!\left(\theta;h_{t-1}\right)$, whereas the prior is only implicitly defined in the data–estimation pairing. In §3.1 we exploited this flexibility through our definition of $p_{f}(\theta;h_{t-1})$. On the other hand, the data–estimation pairing could provide some beneficial flexibility in specifying how the data itself is simulated through changing $p\!\left(y_{t};[h_{t-1},x_{t}]\right)$, which could, for example, be useful when we have access to external data simulators.
@@
This difference becomes particularly noticeable when the complexities of the spaces of $\theta$ and $y$ differ significantly. Our ability to draw sensible samples of either will generally be quite robust to these spaces being complex or high-dimensional; this is where LLMs tend to thrive, effectively generating highly complex outputs in an autoregressive manner. However, evaluating the entropy of a distribution becomes dramatically harder as the dimensionality or complexity increases (Acharya et al., 2019; Paninski, 2003), and the entropy of the predictive distribution of an LLM in such cases will not typically provide a sensible measure of uncertainty (Kadavath et al., 2022; Desai & Durrett, 2020). As such, the decision between joint formulations should predominantly be based on the complexity of the space of $\theta$ versus that of $y$: we should generally favor the prior–likelihood formulation if $\theta$ is more complex and the data–estimation formulation if $y$ is more complex. For the problems that we consider, the space of $y$ is less complex than that of $\theta$, indicating we should, in general, use the prior–likelihood formulation. However, in cases where this is not true, the data–estimation formulation may be preferable instead, see Appendix˜B for further discussion on this and on the choice of $\theta$.
@@ 5 Related work
Several works have explored the baseline ability of LLMs to rapidly learn about a parameter of interest by asking questions (Zhang et al., 2024; Li et al., 2025b)—effectively our Prompt-Only baseline in §6. While these works demonstrate some ability to adaptively construct information-seeking questions, they often fail to extract important information (Li et al., 2025a).
Some works have further specifically attempted to choose questions based on model-based uncertainty criteria (Piriyakulkij et al., 2023; Hu et al., 2024; Kobalczyk et al., 2025; Mazzaccara et al., 2024; Cooper et al., 2025). None of these works provide the same careful consideration of how the underlying joint model should be formulated, which underpins our own work, and they all assume deterministic likelihood models that mean their objectives correspond to a sample-based estimate of marginal predictive entropy in practice, as explained in Section˜3.3. In general, these previous works have also required restrictions on the space of allowable hypotheses, $\theta$, and typically require additional assumptions and/or approximations. More extensive discussion of related work is given in Appendix˜C.
@@ 6 Experiments
We produce answers to the questioner LLM’s questions using a separate answerer LLM. The answerer is provided with a ground-truth $\theta^{*}$ (a target entity in 20 Questions or a user profile in preference elicitation) and processes individual questions from the questioner without access to any of the questioner’s context (i.e. $h_{t-1}$ and the questioner’s prompts). We test two questioner–answerer setups, where the two are served by separate instances of the same LLM, or two different LLMs. The latter scenario is important because in practice, the answerer will often follow a different distribution than the questioner’s internal model for reasoning about responses, thereby forming a model misspecification.
@@
We compare BED-LLM against three baselines. Prompt-Only prompts the questioner to directly generate an informative next question, without explicit hypothesis generation or a data-acquisition objective, and then sampling the question with temperature T=1T\!=\!1; this was explored by Zhang et al. (2024). CoT augments Prompt-Only with a ReAct-style (Yao et al., 2023b) chain-of-thought: the LLM first produces a thought (what has been established, what information would help most) then an action (the question). This tests whether structured reasoning alone can close the gap to BED-LLM. Split chooses the question that most equally splits a sampled set of hypotheses, $\Theta^{\mathrm{cand}}$, which corresponds to maximizing the marginal predictive entropy, $\mathrm{H}[p(y_{t};x_{t},h_{t-1})]$, in a model with a deterministic likelihood. As such, the methods of Cooper et al. (2025), Hu et al. (2024), Kobalczyk et al. (2025), Mazzaccara et al. (2024) and Piriyakulkij et al. (2023) can all be viewed as variants of this objective. To the best of our knowledge, Split represented the previous state-of-the-art method for 20 Questions. We note that our own specific Split baseline implementation, which uses BED-LLM’s filtering mechanism, also achieves dramatically better results than reported by, for example, Kobalczyk et al. (2025), so this constitutes a strong baseline relative to previous work. On top of this, our implementation of Prompt-Only appears to significantly improve over that of Zhang et al. (2024). Additional evaluations and full plots are in Appendix˜F.
@@
We consider three sets of 20 Questions problems: Animals, Celebrities, and Things (See Section˜H.1). Each problem set comprises 100 target entities, $\{\theta^{*}_{i}\}_{i=1}^{100}$. The space of possible $\theta$ is large and not explicitly defined or restricted: we do not tell the LLM this set of target entities, so the space is bounded only by what the LLM can generate; by comparison, many previous works have relied on restricted spaces for $\theta$ (Chan et al., 2025; Hu et al., 2024; Piriyakulkij et al., 2023; Wang et al., 2025).
@@
Model Prompt-Only Split CoT BED-LLM Entropy Data–Est. ICL Beliefs Impl. Max.
@@
[…]
Table 1: Success rate (%) for 20 Questions at the end of the game. Best result in bold. $\pm$ numbers show the standard error of the mean estimated using $\sqrt{p(1-p)/(n-1)}$ where $p$ is the success percentage and $n$ is the number of datapoints. This estimator is positively biased and thus conservative.
Method Joint model Objective Belief updates
BED-LLM Prior–likelihood Full EIG (Eq. 3) Filtered ($p_{f}$)
[…]
Table 2: Summary of how each method relates to BED-LLM’s three core algorithmic components. Each ablation (bottom section) modifies exactly one component of BED-LLM, highlighted in grey.
[…]
BED-LLM improves over Prompt-Only and Split baselines
Our results in Figure˜2 and Table˜1 show BED-LLM significantly outperforming both baselines across all problems and LLMs. Particularly notable is that BED-LLM’s final success rate is typically more than double that of Prompt-Only, highlighting the substantial gains that can be achieved by using explicit EIG maximization.
@@
We further evaluate four ablations of BED-LLM to isolate the contribution of each of its core components; full descriptions are in Appendix˜G. Entropy replaces the full EIG objective with the marginal predictive entropy, $\mathrm{H}[p(y_{t};x_{t},h_{t-1})]$; it is similar to the Split baseline, but uses BED-LLM’s likelihood model instead of a deterministic one. Data–Estimation swaps BED-LLM’s prior–likelihood joint model for a data–estimation pairing (see Appendix˜D), testing the importance of BED-LLM measuring uncertainty in $y$ space instead of $\theta$ space (see Section˜3). ICL Beliefs omits our filtering procedure in the belief extraction (see Section˜3.1), just using the, often incoherent, raw in-context beliefs $p_{\mathrm{LLM}}(\theta;h_{t-1})$ instead of $p_{f}(\theta;h_{t-1})$. Finally, Implicit Maximization replaces explicit EIG estimation with LLM judgment, drawing on the Tree of Thoughts (ToT; Yao et al., 2023a) intuition of reasoning over hypothetical futures: the model is presented with the same candidate questions as BED-LLM, but is simply prompted to internally reason and select the question it judges most informative, providing a lightweight alternative to explicit EIG estimation.
In Table˜1 we see that BED-LLM comfortably outperforms all alternative approaches. Notably, Entropy provides the strongest alternative, with performance slightly better than Split. The fact that Entropy performs much more similarly to Split than BED-LLM shows that the use of a non-deterministic likelihood is beneficial predominantly in allowing us to target a proper EIG, rather than due to changes in the marginal predictive entropy itself. The improvement of Implicit Maximization over Prompt-only is also notable, given how cheap it is to run (see Table˜3 for run time information).
@@
Our analysis in §4 is validated by our results: BED-LLM’s prior-likelihood approach substantially outperforms Data–Estimation. Data–Estimation still outperforms Prompt-Only, but interestingly it performs worse than Entropy, highlighting the importance of estimating uncertainty in the $y$ space instead of $\theta$ space here. These findings reinforce our claim that the choice of joint-model factorization is a critical algorithmic decision.
@@
We also see how we produce our beliefs over $\theta$ matters: deriving beliefs using simple in-context learning, as in ICL Beliefs, lead to massive performance drops. Further, while BED-LLM’s routines for sampling candidate questions and hypotheses are crucial, they alone are not sufficient: passing the samples to an LLM and prompting it to select the highest-EIG question, as in Implicit Maximization, works much less well.
@@
Our results in Figure˜2 demonstrate that the benefit of BED-LLM persists even under model misspecification. This is important for applicability to real-world users, whose responses will follow a different distribution than the questioner LLM.